Two particles having mass ratio of 4:5 have same kinetic energy .Calculate the ratio of their linear momentum
1
Expert's answer
2018-10-08T15:23:08-0400
The kinetic energy of a particle
T=(mv^2)/2
The linear momentum of a particle
p=mv
Thus
T=p^2/2m,p=√2mT
The ratio
p_1/p_2 =√(2m_1 T_1 )/√(2m_2 T_2 )=√(m_1/m_2 )=√(4/5)=2/√5=(2√5)/5